OpenStudy (anonymous):

Identify all values of t E (0,2pi] that satisfy sec(t)= -2 square root of 3 all over 3 I need an example to help me solve this

OpenStudy (e.mccormick):

I am going to go home... but here are two paces with examples: Some text www.regentsprep.org/Regents/math/algtrig/ATT10/trigequations.htm A video that does a simplification first, then shows the solutions over a given domain. http://youtu.be/IJ_ljBdllfw

OpenStudy (mertsj):

\[\sec t=\frac{-2\sqrt{3}}{3}\] \[\cos t=-\frac{3}{2\sqrt{3}}\]